# Tunneling in a Lorenz-like model for an active wave-particle entity

### v1

@@ Abstract
Active wave-particle entities (WPEs) emerge as self-propelled oil droplets on the free surface of a vibrating oil bath. The particle (droplet) periodically imprints decaying waves on the liquid surface which in turn guide the particle motion, resulting in a two-way coupling between the particle and its self-generated waves. Such WPEs have been shown to exhibit hydrodynamic analogs of various quantum features. In this work, we theoretically and numerically explore a dynamical analog of tunneling by considering the setup of a one-dimensional WPE incident on an isolated Gaussian potential barrier. Our idealized model takes the form of a perturbed Lorenz system which we use to explore the dynamics and statistics of barrier crossing as a function of initial conditions and system parameters. Our work highlights that velocity fluctuations of the WPE at high memories that are rooted in non-equilibrium features of the Lorenz system, such as spiraling motion towards equilibrium points and transient chaos, give rise to - (i) sensitivity and unpredictability in barrier crossing, (ii) smooth variations in transmission probability as a function of system parameters, and (iii) wave-like features in the transmitted and reflected probability density profiles.
@@ I Introduction
Active particles are non-equilibrium entities that consume energy and convert it into directed motion [1]. They can be living organisms such as bacteria, algae, animals and birds, or inanimate entities such as colloids or robots [2]. They span a large range of scales from micrometer-sized molecular motors [3] to crowds of humans covering hundreds of meters [4]. A curious inanimate hydrodynamic active system is that of walking [5, 6] and superwalking [7] droplets. In this active system, periodically bouncing millimeter-sized droplets walk horizontally on the free surface of a vertically vibrating bath of the same liquid. Upon each impact, the droplet imprints a localized slowly decaying standing wave on the free surface. The droplet then interacts with these self-generated waves on subsequent impacts which guides the droplet motion and give rise to a self-propelled wave-particle entity (WPE). For high amplitudes of vertical bath vibrations, the droplet-generated waves decay very slowly in time and the walking motion of the droplet is governed by the history of waves generated by the droplet along its trajectory, giving rise to path memory in this active system. Such active WPEs have been shown to mimic several peculiar features that are usually associated with the microscopic quantum realm [8].
@@
Since these active WPEs exhibit hydrodynamic quantum analogs, the framework of generalized pilot-wave dynamics [9] has been formulated which allows the exploration of broader class of dynamical systems and parameter-space for active WPEs that are not restricted by experimental constraints. This theoretical framework has motivated exploration of idealized one-dimensional WPE setups to explore quantum analogs and non-equilibrium behaviors [10, 11, 12, 13, 14, 15, 16, 17]. Since the motion of WPEs is driven by path memory, typical models take the form of integro-differential equations of motion  [18, 19]. However, for $1$D active WPEs with certain simple wave forms, the infinite-dimensional system of integro-differential equation converts to a Lorenz-like nonlinear dynamical system comprising of a few ordinary differential equations (ODEs) [10, 11, 14, 15]. For these systems, the dynamics of the active WPE can be rationalized in terms of the nonlinear and chaotic dynamics of Lorenz-like systems [15, 13]. Such reduced-order descriptions allow for rapid and detailed exploration of the parameter space and initial conditions of the system.
@@
One of the quantum analogs exhibited by these active WPEs is that of tunneling [20, 21, 22, 23, 24, 25]. Eddi et al. [20] first showed in experiments that when a walking droplet is incident on a submerged barrier, it can unpredictably cross the barrier. This was further explored in experiments by Tadrist et al. [22] who showed that the unpredictability is due to sensitivity in the vertical bouncing dynamics of the droplet when interacting with the barrier, as opposed to uncertainty in initial conditions or noise in the experiment. The above experiments were done in a regime where a single walker in free space moves steadily at a constant speed. Recent experiments have shown that at high memories, this steady walking state can become unstable and the walking droplet develops velocity oscillations [26]. In this paper, we theoretically and numerically consider a one-dimensional setup of a WPE incident on a Gaussian potential barrier with a particular focus on the effects of unsteady walking dynamics arising in the high memory regime. Our idealized model with the Gaussian barrier only captures the horizontal walking motion in $1$D for the WPE and takes the form of a perturbed Lorenz system. We investigate in detail, the dynamics and statistics of barrier crossing as a function of system parameters and initial conditions.
@@ II Theoretical Model
Figure 1: Schematic of the system showing a one-dimensional active WPE (blue circle) directed at an isolated Gaussian potential barrier (red). The particle located at $x_{d}$ and moving horizontally with velocity $\dot{x}_{d}$ experiences – a propulsion force from its self-generated wave field $h(x,t)$ (blue filled area), an effective drag force and an external force $F_{ex}$ when interacting with the Gaussian barrier. The underlying wave field $h(x,t)$ is a superposition of the individual waves of spatial form $W(x)$ that decay exponentially in time (black and gray curves with the grayness indicating waves generated at earlier times), and are generated by the particle continuously along its trajectory.
@@ IV.2 Dynamics of barrier crossing with increasing memory
Since the WPE dynamics without the potential barrier maps to the Lorenz system [14, 17], we proceed by reviewing different attractors realized in the Lorenz model [29, 17] of Eq. (3) (with $H=0$), as a function of the memory parameter $\tau$. For a fixed $R$ and small values of $\tau$ (i.e. $\tau<1/\sqrt{R}$), there is only one equilibrium point of the system at $(X,Y,Z)=(0,0,0)$ which is stable. This corresponds to a stationary non-walking state for the particle since its velocity is $\dot{x}_{d}=X=0$. With increasing $\tau$, the Lorenz system undergoes a pitchfork bifurcation (at $\tau=1/\sqrt{R}$) where the non-walking state becomes unstable and a symmetric pair of stable equilibrium points emerge, $(X,Y,Z)=(\pm\sqrt{R-1/\tau^{2}},\pm\sqrt{R-1/\tau^{2}},1/\tau)$; this corresponds to steady walking motion of the particle to right/left. Further increase in $\tau$ leads to a complex set of global bifurcations in the Lorenz system [29, 30] and eventually chaos emerges at large $\tau$ with a strange attractor in the phase space. Phase-space dynamics on the Lorenz chaotic attractor corresponds to a chaotic walk [14, 31] for the active particle where it unpredictably switches between left and right moving states as the phase space trajectory jumps between left and right “wing” of the Lorenz attractor.
@@
Further increasing memory parameter to $\tau=1.81$, the steady walking equilibrium points are still stable, however, their eigenvalues become complex conjugates and they acquire the character of stable spirals in phase space. Hence, the approach to these fixed points in phase-space will have an oscillatory component and this corresponds to velocity oscillations of the WPE as it reaches its steady state velocity. Figure 5(a) shows the basin of attraction of different types of trajectories in the initial condition plane. For WPE starting far away from the barrier, the transient effects have decayed and the WPE converges to the steady walking state when it approaches the barrier. Since this steady walking speed is still not sufficient to cross the potential barrier, we have reflected trajectories and a corresponding blue region for $x_{d}(0)<<x_{b}$. For an intermediate range of starting positions away from the potential barrier, we see alternating red and blue stripes in Fig. 5(a). These patterns in the basin of attraction arise from velocity oscillations in the WPE motion as it approaches the barrier. Since the transient velocity oscillations have not decayed by the time the WPE reaches the barrier, The WPE can have a larger or a smaller velocity than the critical velocity required to cross the barrier when it starts interacting with the barrier (see Fig. 5(b)-(c)) and this dictates whether the WPE transmits through the potential barrier or gets reflected from it (see also Movie 3). Again, due to similar mechanisms as shown for Fig. 4, we have a small green region of non-interacting trajectories and red triangular region of transmitted trajectories for large initial positive velocities near the barrier.
@@
As the memory parameter is increased to $\tau=2$ and $\tau=3$, then as shown in Fig. 6(a), we see two qualitative changes. Firstly, the steady state velocity of the WPE is now sufficient to cross the potential barrier and hence, for WPE starting further away from the barrier we now see transmitted trajectories and a corresponding red region. Secondly, the extent of blue and red stripes are greatly reduced since even with velocity oscillations, the minimum velocity is still enough to cross the potential barrier for most part of the initial condition space shown in Fig. 6(a). As the memory parameter is further increased to $\tau=3.6$, as shown in Fig. 6(b)-(c), we see emergent patches of intermingled basins (see the inset of Fig. 6(c)) between the transmitted and reflected states due to the presence of transient chaos in the Lorenz system in this parameter range. This is illustrated by showing the position-time (see Fig. 6(d)) and projection of phase-space (see Fig. 6(e)) of two trajectories with closely spaced initial conditions in the intermingled basin region. These two closely starting WPEs follow almost identical trajectories in phase-space before interacting with the potential barrier. As they both interact and cross the potential barrier, their almost identical phase-space trajectories are perturbed in different ways. One of the perturbed phase-space trajectories (red), quickly settles back on the same side of the spiral fixed point after a short excursion on the other wing of the Lorenz attractor. The other perturbed phase-space trajectory (blue) gets trapped in transient chaos after crossing the potential barrier. This transient chaos results in chaotic walking for this WPE and it interacts with the potential barrier again and crosses it. During this second interaction, the phase-space trajectory is again perturbed and settles towards the negative velocity spiral fixed point and the WPE ends up on the same side of the potential barrier as the incident trajectory (see also Movie 4). Hence we see a final state sensitivity [32] between transmitted and reflected states due to the presence of transient chaos at high memories for the WPE.
@@ V Statistics of barrier crossing
After having explored the different dynamical mechanisms for barrier crossing in the previous section, we now turn to investigate the statistical properties for barrier crossing. We do this by calculating the transmission probability $\text{Pr}(T)$. This probability is calculated by ensemble averaging over a subset of possible initial conditions. Specifically, we consider the plane of initial conditions where $x_{d}(0)\in[30,70]$ and $X(0)\in[0,3\sqrt{R-1/\tau^{2}}]$ with fixed $Y(0)=Z(0)=0$. This subset of initial conditions corresponds to no initial wave field for the WPE and the initial location of the WPE is far from the Gaussian potential centered at $x_{b}=80$. Further, the WPE initial velocity is in a reasonable range of its steady velocity. These choice of initial conditions would be reasonable for experiments and they also ensure that the crossing statistics are not significantly distorted by rapidly decaying transient effects from initial conditions. The transmission probability is then defined as the fraction of area occupied by red region in the initial condition space relative to the total area covered by blue and red regions (i.e. we discard the green region).
@@ V.3 Evolution of probability density profiles
For a quantum particle interacting with a finite potential barrier, one usually plots the evolution of the probability density profile as a function of time which clearly shows the transmitted and reflected part of the probability density after the initial distribution interacts with the barrier [34]. We create similar space-time plots showing the evolution of a distribution of initial conditions. The initial velocities of all WPEs is set to the steady-state positive velocity i.e. $X(0)=\sqrt{R-1/\tau^{2}}$ while $Y(0)=Z(0)=0$. The initial positions are sampled from a normal distribution with a mean of $60$ and a standard deviation of $5$. This results in an initial normal distribution of initial conditions in position and we plot its evolution in Fig. 9. In Fig. 9(a), for $\tau=1.5$, the probability density profile gets reflected from the potential barrier because the steady-state velocity is insufficient to cross the barrier. In Fig. 9(b), for $\tau=1.81$, even though the initial velocity is set at the steady state velocity, since $Y(0)$ and $Z(0)$ are not at its equilibrium values, we get velocity oscillations. Hence, when the probability density profile interacts with the barrier, part of it is transmitted and part of it is reflected due to the dynamical mechanism noted for Fig. 5. Further, the reflected and transmitted probability density profiles develop multiple peaks. Similar features were observed in the density evolution of a Bose-Einstein condensate (BEC) tunneling through a Gaussian potential barrier where both classical and quantum effects are present [35]. Further increasing memory to $\tau=2$, the WPE velocity is sufficient for crossing the barrier (see Fig. 6(a)) and hence we see most of the probability density profile transmitting through the barrier. However, due to the presence of velocity oscillations, a qualitative change with multiple peaks is observed in the probability distribution after the transmission. Lastly, for $\tau=3.6$ in Fig. 9(d), velocity oscillations are evident even before encountering the potential barrier, and after interaction with the barrier, transient chaos can occur as noted in Fig. 6(c)-(e). A part of the transmitted and reflected probability density profile corresponds to trajectories that experience little or no transient chaos, however, multiple peaks arise due to velocity oscillations. For the initial conditions undergoing transient chaos, they spend a long time in the vicinity of the barrier and then eventually get transmitted or reflected resulting in smaller probability density peaks transmitting or reflecting at later times. We note that in a recent work, Rahman [36] showed transmission and reflection of a probability density profile interacting with a rectangular potential barrier using their discrete-time model of walking droplets.

### v2

@@ Abstract
Active wave-particle entities (WPEs) emerge as self-propelled oil droplets on the free surface of a vibrating oil bath. The particle (droplet) periodically imprints decaying waves on the liquid surface which in turn guide the particle motion, resulting in a two-way coupling between the particle and its self-generated waves. Such WPEs have been shown to exhibit hydrodynamic analogs of various quantum features. In this work, we theoretically and numerically explore a dynamical analog of tunneling by considering a simple setup of a one-dimensional WPE incident on an isolated Gaussian potential barrier. Our idealized model takes the form of a perturbed Lorenz system which we use to explore the dynamics and statistics of barrier crossing as a function of initial conditions and system parameters. Our work highlights that velocity fluctuations of the WPE at high memories that are rooted in non-equilibrium features of the Lorenz system, such as spiraling motion towards equilibrium points and transient chaos, give rise to - (i) sensitivity and unpredictability in barrier crossing, (ii) smooth variations in transmission probability as a function of system parameters, and (iii) wave-like features in the transmitted and reflected probability density profiles.
@@ I Introduction
Active particles are non-equilibrium entities that consume energy and convert it into directed motion [1]. They can be living organisms such as bacteria, algae, animals and birds, or inanimate entities such as colloids or robots [2]. They span a large range of scales from micrometer-sized molecular motors [3] to crowds of humans covering hundreds of meters [4]. A curious inanimate hydrodynamic active system is that of walking [5, 6] and superwalking [7] droplets. In this active system, periodically bouncing millimeter-sized droplets walk horizontally on the free surface of a vertically vibrating bath of the same liquid. Upon each impact, the droplet imprints a localized slowly decaying standing wave on the free surface. The droplet then interacts with these self-generated waves on subsequent impacts which guide the droplet motion and give rise to a self-propelled wave-particle entity (WPE). For high amplitudes of vertical bath vibrations, the droplet-generated waves decay very slowly in time, and the walking motion of the droplet is governed by the history of waves generated by the droplet along its trajectory, giving rise to path memory in this active system. Such active WPEs have been shown to mimic several peculiar features that are usually associated with the microscopic quantum realm [8].
@@
Encouraged by the several hydrodynamic quantum analogs exhibited by these active WPEs, the framework of generalized pilot-wave dynamics [9] has been formulated which allows the exploration of a wider class of dynamical systems and parameter-space that are not restricted by experimental constraints. In particular, this theoretical framework has motivated exploration of idealized one-dimensional pilot-wave models to explore quantum analogs and non-equilibrium behaviors [10, 11, 12, 13, 14, 15, 16, 17, 18]. Since the motion of WPEs is driven by path memory, typical models take the form of integro-differential equations of motion  [19, 20]. However, for $1$D active WPEs with certain simple wave forms, the infinite-dimensional system of integro-differential equation converts to a Lorenz-like nonlinear dynamical system comprising of a few ordinary differential equations (ODEs) [10, 11, 14, 15]. For these systems, the dynamics of the active WPE can be rationalized in terms of the nonlinear and chaotic dynamics of Lorenz-like systems [11, 15, 18]. Such reduced-order ODE models allow for rapid and detailed exploration of the parameter space and initial conditions of the system.
@@
One of the quantum analogs exhibited by these active WPEs is that of tunneling [21, 22, 23, 24, 25, 26]. Eddi et al. [21] first showed in experiments that when a walking droplet is incident on a submerged barrier, it can unpredictably cross the barrier. This was further explored in experiments by Tadrist et al. [23] who showed that the unpredictability is due to sensitivity in the vertical bouncing dynamics of the droplet when interacting with the barrier, as opposed to uncertainty in initial conditions or noise in the experiment. The above experiments were done in a regime where a single walker in free space moves steadily at a constant speed. Recent experiments have shown that at high memories, this steady walking state can become unstable and the walking droplet develops velocity oscillations [27]. In this paper, we theoretically and numerically consider a one-dimensional setup of a WPE incident on a Gaussian potential barrier with a particular focus on the effects of unsteady walking dynamics arising in the high-memory regime. Our aim is to explore in detail the parameter space of the system as well as the space of initial conditions. A complete model that captures the vertical bouncing dynamics, horizontal walking dynamics as well as the wave evolution, would be computationally very expensive to perform such a detailed exploration. Hence, we resort to an idealized Lorenz-like ODE model to investigate in detail the dynamics and statistics of barrier crossing as a function of system parameters and initial conditions. Nevertheless, we expect that the interesting dynamical behaviors identified from the detailed exploration of our simplified model will guide in selecting parameter regimes to probe for future experimental work and numerical exploration with more accurate models. Furthermore, independent of the experimental system of walking droplets, our work demonstrates that a plethora of rich behaviors associated with barrier crossing can arise from the underlying nonlinear and chaotic dynamics of a simple pilot-wave model.
@@ II Theoretical Model
Figure 1: Schematic of the system showing a one-dimensional active WPE (blue) directed at an isolated Gaussian potential barrier (red). The particle (blue circle) located at $x_{d}$ and moving horizontally with velocity $\dot{x}_{d}$ experiences – a propulsion force from its self-generated wave field $h(x,t)$ (blue filled area), an effective drag force and an external force $F_{ex}$ when interacting with the Gaussian barrier. The underlying wave field $h(x,t)$ is a superposition of the individual waves of spatial form $W(x)$ that decay exponentially in time (black and gray curves with the grayness indicating waves generated at earlier times), and are generated by the particle continuously along its trajectory.
@@ IV.2 Dynamics of barrier crossing with increasing memory
Since the WPE dynamics without the potential barrier maps to the Lorenz system [14, 17], we proceed by reviewing different attractors realized in the Lorenz model [30, 17] of Eq. (3) (with $H=0$), as a function of the memory parameter $\tau$. For a fixed $R$ and small values of $\tau$ (i.e. $\tau<1/\sqrt{R}$), there is only one equilibrium point of the system at $(X,Y,Z)=(0,0,0)$ which is stable. This corresponds to a stationary non-walking state for the particle since its velocity is $\dot{x}_{d}=X=0$. With increasing $\tau$, the Lorenz system undergoes a pitchfork bifurcation (at $\tau=1/\sqrt{R}$) where the non-walking state becomes unstable and a symmetric pair of stable equilibrium points emerge, $(X,Y,Z)=(\pm\sqrt{R-1/\tau^{2}},\pm\sqrt{R-1/\tau^{2}},1/\tau)$; this corresponds to steady walking motion of the particle to right/left. Further increase in $\tau$ leads to a complex set of global bifurcations in the Lorenz system [30, 31] that lead to oscillatory dynamics of $X$ and eventually chaos emerges at large $\tau$ with a strange attractor in the phase space. Phase-space dynamics on the Lorenz chaotic attractor corresponds to a chaotic walk [14, 32] for the active particle where it unpredictably switches between left and right moving states as the phase space trajectory jumps between left and right “wing” of the Lorenz attractor. The stationary non-walking state as well as the steady walking states of the WPE are routinely observed in experiments with walking and superwalking droplets moving in free space [6, 7]. Instability of the steady walking state and the presence of velocity oscillations has also been reported for a free walker [27], whereas chaotic walking states at high memory of a WPE moving in free-space have not been realized experimentally yet.
@@
Further increasing memory parameter to $\tau=1.81$, the steady walking equilibrium points are still stable, however, their eigenvalues have become complex conjugates and they acquire the character of stable spirals in phase space. Hence, the approach to these fixed points in phase-space will have an oscillatory component and this corresponds to velocity oscillations of the WPE as it reaches its steady state velocity. Figure 5(a) shows the basin of attraction of different types of trajectories in the initial condition plane. For WPE starting far away from the barrier, the transient effects have decayed and the WPE converges to the steady walking state when it approaches the barrier. Since this steady walking speed is still not sufficient to cross the potential barrier, we have reflected trajectories and a corresponding blue region for $x_{d}(0)<<x_{b}$. For an intermediate range of starting positions away from the potential barrier, we see alternating red and blue stripes in Fig. 5(a). These patterns in the basin of attraction arise from velocity oscillations in the WPE motion as it approaches the barrier. Since the transient velocity oscillations have not decayed by the time the WPE reaches the barrier, the WPE can have a larger or a smaller velocity than the critical velocity required to cross the barrier when it starts interacting with the barrier (see Fig. 5(b)-(c)) and this dictates whether the WPE transmits through the potential barrier or gets reflected from it (see also Movie 3). Again, due to similar mechanisms as shown for Fig. 4, we have a small green region of non-interacting trajectories and red triangular region of transmitted trajectories for large initial positive velocities near the barrier.
@@
As the memory parameter is increased to $\tau=2$ and $\tau=3$, then as shown in Figs. 6(a)-(b), we see two qualitative changes. Firstly, the steady state velocity of the WPE is now sufficient to cross the potential barrier and hence, for WPE starting further away from the barrier we now see transmitted trajectories and a corresponding red region. Secondly, the extent of blue and red stripes are greatly reduced since even with velocity oscillations, the minimum velocity is still enough to cross the potential barrier for most part of the initial condition space shown in Figs. 6(a)-(b). As the memory parameter is further increased to $\tau=3.6$, as shown in Fig. 6(c), we see emergent patches of intermingled basins (see the inset of Fig. 6(c)) between the transmitted and reflected states due to the presence of transient chaos in the Lorenz system in this parameter range. This is illustrated by showing the position-time (see Fig. 6(d)) and projection of phase-space (see Fig. 6(e)) of two trajectories with closely spaced initial conditions in the intermingled basin region. These two closely starting WPEs follow almost identical trajectories in phase-space before interacting with the potential barrier. As they both interact and cross the potential barrier, their almost identical phase-space trajectories are perturbed in different ways. One of the perturbed phase-space trajectories (red), quickly settles back on the same side of the spiral fixed point after a short excursion on the other wing of the Lorenz attractor. The other perturbed phase-space trajectory (blue) gets trapped in transient chaos after crossing the potential barrier. This transient chaos results in chaotic walking for this WPE and it interacts with the potential barrier again and crosses it. During this second interaction, the phase-space trajectory is again perturbed and settles towards the negative velocity spiral fixed point and the WPE ends up on the same side of the potential barrier as the incident trajectory (see also Movie 4). Hence we see a final state sensitivity [33] between transmitted and reflected states due to the presence of transient chaos at high memories for the WPE.
@@ V Statistics of barrier crossing
After having explored the different dynamical mechanisms for barrier crossing in the previous section, we now turn to investigate the statistical properties for barrier crossing. We do this by calculating the transmission probability $\text{Pr}(T)$. This probability is calculated by ensemble averaging over a subset of possible initial conditions. Specifically, we consider the plane of initial conditions where $x_{d}(0)\in[30,70]$ and $X(0)\in[0,3\sqrt{R-1/\tau^{2}}]$ with fixed $Y(0)=Z(0)=0$. This subset of initial conditions corresponds to no initial wave field for the WPE and the initial location of the WPE is far from the Gaussian potential centered at $x_{b}=80$. Further, the WPE initial velocity is in a reasonable range of its steady velocity. These choice of initial conditions would be reasonable for the experimental walking droplet system, and they also ensure that the crossing statistics are not significantly distorted by rapidly decaying transient effects from initial conditions. The transmission probability is then defined as the fraction of area occupied by red region in the initial condition space relative to the total area covered by blue and red regions (i.e. we discard the green region).
@@ V.3 Evolution of probability density profiles
For a quantum particle interacting with a finite potential barrier, one usually plots the evolution of the probability density profile as a function of time which clearly shows the transmitted and reflected part of the probability density after the initial distribution interacts with the barrier [35]. We create similar space-time plots showing the evolution of a distribution of initial conditions. The initial velocities of all WPEs is set to the steady-state positive velocity i.e. $X(0)=\sqrt{R-1/\tau^{2}}$ while $Y(0)=Z(0)=0$. The initial positions are sampled from a normal distribution with a mean of $60$ and a standard deviation of $5$. This results in an initial normal distribution of initial conditions in position and we plot its evolution in Fig. 9. In Fig. 9(a), for $\tau=1.5$, the probability density profile gets reflected from the potential barrier because the steady-state velocity is insufficient to cross the barrier. In Fig. 9(b), for $\tau=1.81$, even though the initial velocity is set at the steady state velocity, since $Y(0)$ and $Z(0)$ are not at its equilibrium values, we get velocity oscillations. Hence, when the probability density profile interacts with the barrier, part of it is transmitted and part of it is reflected due to the dynamical mechanism noted for Fig. 5. Further, the reflected and transmitted probability density profiles develop multiple peaks. We note that qualitative similar features of multiple peaks in the probability density have been observed in numerically simulated Bose-Einstein condensate (BEC) tunneling through a Gaussian potential barrier where both classical and quantum effects are present [36]. Further increasing memory to $\tau=2$, the WPE velocity is sufficient for crossing the barrier (see Fig. 6(a)) and hence we see most of the probability density profile transmitting through the barrier. However, due to the presence of velocity oscillations, a qualitative change with multiple peaks is observed in the probability distribution after the transmission. Lastly, for $\tau=3.6$ in Fig. 9(d), velocity oscillations are evident even before encountering the potential barrier, and after interaction with the barrier, transient chaos can occur as noted in Fig. 6(c)-(e). A part of the transmitted and reflected probability density profile corresponds to trajectories that experience little or no transient chaos, however, multiple peaks arise due to velocity oscillations. For the initial conditions undergoing transient chaos, they spend a long time in the vicinity of the barrier and then eventually get transmitted or reflected resulting in smaller probability density peaks transmitting or reflecting at later times. We note that in a recent work, Rahman [37] showed transmission and reflection of a probability density profile interacting with a rectangular potential barrier using their discrete-time model of walking droplets.